Circular No. 7393

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

CYGNUS X-3 D. Dal Fiume, E. Palazzi, and M. Orlandini, ITESRE, CNR, Bologna;

and L. Nicastro and S. Del Sordo, IFCAI, CNR, Palermo, on behalf of a large collaboration, write: “BeppoSAX observed Cyg X-3 on Mar. 21.779– 22.599 UT for ∼ 30 000 s during an extended radio and hard-x-ray quenched state (IAUC 7365). The source is well detected in the PDS instrument up to 140 keV at a flux level of ∼ 10 mCrab. A quick analysis shows major and complex changes in the emitted x-ray spectrum when compared to the normal quiescent radio state. Measured source fluxes in given energy bands: 7 × 10^{−9} erg s^{−1} cm^{−2}, 2–10 keV; 4.6 × 10^{−10}, 10–20; 1.2 × 10^{−10}, 20–100. A 10-percent systematic error in these fluxes must be added due to a rather poor spectral fit. No model that we previously used gave a satisfactory fit. For comparison, the fluxes measured during the quiescent radio state in Sept. 1997 in the same energy bands were respectively 2.9 × 10^{−9}, 1.8 × 10^{−9}, and 1.7 × 10^{−9} erg s^{−1} cm^{−2}. Additional BeppoSAX observations are planned.”

BC URSAE MAJORIS E. O. Waagen, AAVSO, reports that this SU UMa-type cataclysmic

variable is in outburst for the first time since July 1995, according to observations in the AAVSO International Database, with the following recent visual magnitude estimates showing that the present event is likely a superoutburst: Mar. 26.870 UT, [14.0 (E. Muyllaert, Oostende, Belgium); 27.256, [15.3 (G. Hanson, Cave Creek, AZ); 28.80, [13.7 (P. Schmeer, Bischmisheim, Germany); 31.799, 12.1 (Schmeer); 31.836, 11.3 (T. Kinnunen, Espoo, Finland); Apr. 1.06, 11.0 (G. Dyck, Assonet, MA); 2.139, 11.6 (Muyllaert); 3.255, 12.1 (Hanson); 4.013, 13.1 (Schmeer).

COMET C/2000 F1 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Center,

reports observations (measured by D. Hammer and himself, reduced by B. G. Marsden) of another Kreutz sungrazing comet found by M. Meyer and J. D. Shanklin via the SOHO website. The comet was observed in both the LASCO C2 and C3 telescopes and showed clear evidence for a tail. Astrometry and orbital elements appear on MPEC 2000-G18.

2000 UT  α_{2000}  δ_{2000}
Mar. 31.846  0^{h}55.^{m}4  + 3^{o}34′
2000 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green